OpenStudy (anonymous):

Solve the proportion. Round to four significant digits if necessary. 5/7=x-6/x+5

OpenStudy (rizags):

could you type it using the equation editor because im not sure which terms are in the numerator

OpenStudy (rizags):

is it like this? \[\frac{ 5 }{ 7 }=\frac{ x-6 }{x+5 }\]

OpenStudy (anonymous):

yes

OpenStudy (rizags):

cross multiply them. do you know how to do that?

OpenStudy (calculusxy):

Cross multiplying is just basically:  From there you will get two equivalent equations: \[7(x-6)=5(x+5)\]